How many litres of pure antifreeze must be added to 6 litres of a 10% antifreeze solution to obtain a 20% anitfreeze solution.
6L=10% so does 12L=20% ???

Answer by stanbon(75887)

How many litres of pure antifreeze must be added to 6 litres of a 10% antifreeze solution to obtain a 20% anitfreeze solution.
-------
Equation:
anti + anti = anti
0.10*6 + 1*x = 0.20(6+x)
-------
10*6 + 100x = 20*6 + 20x
------
80x = 10*6
x = (1/8)6
x = 0.375 litres (amt of pure antifreeze to add)
===============
Cheers,
Stan H.
===============